# TetSphere Splatting: Representing High-Quality Geometry with Lagrangian Volumetric Meshes

### v2

@@ 6.2 Results
Single-view Reconstruction. We use the $8$ colored images and normal maps generated by Wonder3d to reconstruct the shapes. Table 1 shows the comparison results. Our TetSphere splatting technique excels over baseline methods in terms of mesh quality while also achieving competitive levels of reconstruction accuracy (Fig. 1(b)). This improvement is attributed to the regularizations embedded within the TetSphere splatting optimization. These results demonstrate the superior capabilities of our representation compared to existing ones.
Image-to-3D Shape Generation. Fig. 6 and 7 illustrate the comparison results on image-to-3D shape generation. Our approach outperforms Magic123 and DreamCraft3D regarding mesh quality, achieving smoother surfaces for broad regions, such as animal bodies while retaining local sharp details in areas like eyes and noses. Magic123 tends to produce overly smooth surfaces that sometimes deviate from the correct geometry, as observed in the mushroom and the teddy bear in Fig. 7. DreamCraft3D suffers from noisy and bumpy surface meshes, indicating lower mesh quality. Furthermore, we highlight the computational efficiency of TetSphere splatting in Table 2. Compared to various geometry representations, TetSphere splatting stands out for its minimal memory usage and achieves the fastest run-time speed. This efficiency underscores the benefits of TetSphere’s explicit and Lagrangian properties.
Figure 9: Results on text-to-3D shape generation. Our TetSphere splatting excels in creating slender forms, demonstrating its strength in handling thin structures effectively.
Text-to-3D Shape Generation. Fig. 9 shows comparison results on text-to-3D generation, with additional results provided in Appendix 0.C. Our TetSphere splatting is capable of producing slender structures such as the dragon’s head and the goblet. Furthermore, the results demonstrate that the integration of TetSphere splatting with SDS enhances the geometric detail of the generated shapes, resulting in both diverse and high-quality textures.
@@ 6.3 Analysis
Effects of Energy Coefficients. Fig. 8 demonstrates how different energy coefficients influence the reconstruction outcome. Larger coefficients lead to a smoother surface, but excessively small coefficients may cause tetrahedron inversion. In our experiments, we choose $w_{1}=5\times 10^{-6},w_{2}=2\times 10^{-5}$ and apply a cosine increasing schedule to balance surface smoothness and structural integrity.
Dense-view Inverse Rendering. In Appendix 0.B, we show an additional application of TetSphere splatting: inverse rendering with densly sampled views and compare it with surface-mesh representation [65]. Our method demonstrates fast and robust optimization results in this context. This further illustrates TetSphere splatting’s potential for broader reconstruction scenarios.
@@ 7 Conclusion
We introduced TetSphere splatting, a geometry representation for reconstructing textured shapes using a tetrahedral mesh framework. This method addresses the limitations of existing reconstruction methods, such as the high computational cost of neural implicit representations and the suboptimal mesh quality inherent in Eulerian geometry methods. Future work could extend TetSphere splatting to leverage direct 3D supervision with volumetric data. The current limitation of TetSphere splatting lies in its inability to guarantee topology preservation due to the union of all tetrahedral spheres. This underscores the necessity for future development of shape generation that can adhere to topology constraints.
Appendix 0.A Implementation Details
Baselines. For the baseline methods, including RealFusion [56], Magic123 [68], One-2-3-45 [47], Point-E [64], Shap-E [39], Zero123 [49], SyncDreamer [50], Wonder3d [53], Open-LRM [33, 31], DreamGaussian [83], Dreamcraft3D [80], and RichDreamer [69], we follow their original implementations and use their publicly available codebase for getting the results in this paper.
Our method. Our implementation of the TetSphere initialization algorithm is developed in C++ and makes use of the Gurobi linear programming solver. The optimization of geometric energies is implemented using CUDA as a PyTorch extension to enhance computational efficiency. 1) For single-view reconstruction, we employ Wonder3d [53] to generate six multi-view images, including both color and normal images for each view, using predefined camera poses for reconstruction with TetSphere splatting. We use a 2-layer MLP for texture representation. The optimization objective $\mathbf{\Phi}(\cdot)$ encompasses both the rendering loss and the normal loss. The rendering loss consists of an $l_{1}$ norm on tone-mapped color and MSE on the alpha mask, along with a cosine loss on normals, similar to that described in [63]; 2) For image-to-3D shape generation, we obtain multi-view images from the initial stage of DreamCraft3D (coarse NeRF fitting only), generating $360$ views sampled on a Fibonacci sphere [28]. The texture optimization directly employs a $2048\times 2048$ 2D texture image, circumventing the need for additional neural networks. Here, the optimization objective $\mathbf{\Phi}(\cdot)$ focuses solely on the rendering loss; 3) For text-to-3D generation, we use the initial stage of RichDreamer [69] to obtain multi-view images. In this scenario, we optimize our TetSphere splatting using both the rendering loss and the SDS loss. The SDS loss for geometry is calculated using the Normal-depth diffusion model as described in [69]. Likewise, the SDS loss for texture leverages the albedo diffusion model from the same work. We implement a 2-layer MLP to parameterize and optimize the PBR material of the shape. To enhance robustness and efficiency, we apply a cosine scheduler to scale the coefficients of the geometry loss for all applications, formulated as $\eta=4^{\sin(\frac{t\pi}{2n})}$, where $t$ denotes the current iteration and $T$ represents the total number of iterations.
Appendix 0.B Dense-view Inverse Rendering
Figure 10: Results on 3D shape reconstruction from dense multi-view images (1/2). Shapes reconstructed using our method are depicted in gray, whereas those reconstructed with the state-of-the-art method from [65] are presented in blue. While both methods effectively minimize the rendering loss, our technique does not produce artifacts, such as undesired wrinkles and flipping triangles (highlighted in red), and producing high-quality triangles, as illustrated by the wireframes.
Fig. 10, 11, and 12 show the results of 3D shape reconstruction from dense multi-view images. These multi-view images were captured by densely sampling camera positions around each object (employing 360 views sampled on a sphere based on the Fibonacci sequence [28]) and subsequently rendered using Mitsuba [36]. The chosen shapes for this demonstration are from the GSO dataset.
To establish the efficacy of our method, we compare its performance with the state-of-the-art surface mesh-based approach as described in [65] by applying both techniques to the aforementioned dense multi-view images. Our method utilizes a consistent hyperparameter set for all shapes undergoing $3,000$ optimization iterations, mirroring the parameters described in the main paper. In contrast, for the method in [65], we adhere to the parameters in the original paper – namely, using $\lambda=19$ and a step size of $10^{-2}$ – and conduct $10,000$ iterations of optimization. This number of iterations is necessary due to difficulties in achieving convergence within a few thousand iterations for this method.
While both methods effectively minimize the rendering loss, our method consistently delivers results of superior quality. Fig. 10 and 11 illustrate that our reconstructions are devoid of common artifacts such as wiggles or kinks, and showcase high-quality triangles, as evidenced by the wireframe representations. Furthermore, our approach demonstrates faster convergence and notable scalability advantages when compared to the second-order method in [65]. Additionally, as Fig. 12 indicates, our method is also capable of handling shapes with complex topologies, further underscoring its versatility and effectiveness.
Appendix 0.C Additional Results
0.C.1 Qualitative Results on Single-view Reconstruction
@@
Fig. 13 shows the results of single-view reconstruction performed on the GSO dataset. Our TetSphere demonstrates superior mesh quality, effectively capturing sharp geometric features, including the boundaries of the shoes.
Figure 11: Results on 3D shape reconstruction from dense multi-view images (2/2).
Figure 12: Our method is capable of handling shapes with complex topologies, an advantage that existing methods fail to achieve [65]. In this example, both shapes are computed by incorporating the rendering loss and normal loss.
Figure 13: Qualitative results on single-view reconstruction.
0.C.2 More Results on Text-to-3D Shape Generation
@@
Figure 14: More results on text-to-3D shape generation.
Appendix 0.D Tetrahedron and its Deformation Gradient
Following [79], we treat a tetrahedron as a piecewise linear element. The initial (undeformed) positions of the four vertices of a tetrahedron are denoted by $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\mathbf{X}^{(3)},\mathbf{X}^{(4)}]$, with each $\mathbf{X}^{(i)}\in\mathbb{R}^{3}$. Similarly, the positions of the four vertices of a deformed tetrahedron are represented by $\mathbf{x}=[\mathbf{x}^{(1)},\mathbf{x}^{(2)},\mathbf{x}^{(3)},\mathbf{x}^{(4)}]$, where each $\mathbf{x}^{(i)}\in\mathbb{R}^{3}$. The deformation gradient $\mathbf{F}\in\mathbb{R}^{3\times 3}$, which quantifies the local deformation of the tetrahedron, is given by:
@@ Appendix 0.D Tetrahedron and its Deformation Gradient
Appendix 0.E Formulation Details of TetSphere Initialization
@@
Appendix 0.F Related work – Text-to-3D content generation

### v4

@@ 5.3 Results
Multi-view reconstruction. Table 1 and Table 5 in Appendix E.1 present the quantitative comparison results. Fig. 5 shows the qualitative comparison on four different shapes. Our method excels over baseline methods regarding mesh quality while achieving similar reconstruction accuracy. Specifically, our approach features uniform meshing (as indicated by ALR) and eliminates floating artifacts (as shown by CC Diff.). This improvement is attributed to the volumetric representation of TetSphere and its geometric energy regularization. The qualitative results further support these findings: our method accurately captures slender and thin structures, such as the fingers of Lucy’s statue, the sorter, and the dress, which other methods struggle with. Additionally, our method demonstrates correct surface orientation, whereas baseline methods, particularly FlexiCubes, and DMesh, often exhibit inverted triangles, as seen in the noisy rendered normal images. However, our TetSphere approach is less effective when modeling thin shells, such as the leaves of the plant, where DMesh shows better accuracy.
Single-view reconstruction. Table 2 shows the comparison results. Fig. 11 in Appendix E.2 shows qualitative results of our method. Our TetSphere splatting excels over baseline methods in terms of mesh quality while also achieving competitive levels of reconstruction accuracy. This improvement is attributed to the regularizations embedded within the TetSphere splatting optimization, ensuring uniform meshing and eliminating common artifacts like floating components. These findings are consistent with the results observed in multi-view reconstruction.
Figure 7: Results on text-to-3D shape generation. Our TetSphere splatting excels in handling slender, thin structures, such as the wine glass stem and the dragon head. More results are available in Appendix E.4.
@@ 5.4 Analysis
Applications to generative modeling. TetSphere splatting is a versatile approach that can be effectively applied to generative modeling tasks. We showcase two applications: image-to-3D and text-to-3D generation. Fig. 12 in Appendix E.3 illustrate the results on image-to-3D generation. Our approach outperforms Magic123 and DreamCraft3D regarding mesh quality, achieving smoother surfaces for broad regions, such as animal bodies while retaining local sharp details in areas like eyes and noses. Magic123 tends to produce overly smooth surfaces that sometimes deviate from the correct geometry. DreamCraft3D suffers from noisy and bumpy surface meshes, indicating lower mesh quality. Fig. 7 and Fig. 13 in Appendix E.4 shows comparison results on text-to-3D generation. Our TetSphere splatting can produce slender structures such as the dragon’s head and the goblet. Furthermore, the results demonstrate that using TetSphere splatting with SDS enhances the geometric detail of the generated shapes, resulting in both diverse and high-quality textures.
Analysis on computational cost As shown in Table 5 in Appendix E.1, our method outperforms the Eulerian method NIE regarding running time while remaining competitive with other geometry representation methods. To further assess the performance of TetSphere splatting, we conduct an extreme test by applying it to the task of image-to-3D generation with SDS loss. We evaluate several baseline methods and report the maximum batch size of $256\times 256$ images that can be processed on a 40GB A100 GPU, as well as the training run-time speed with a batch size of $4$, as shown in Table 3. TetSphere splatting stands out for its minimal memory usage and achieves the fastest run-time speed. This efficiency underscores the benefits of TetSphere’s explicit and Lagrangian properties.
Effects of energy coefficients. Fig. 6 demonstrates how different energy coefficients influence the reconstruction outcome. Larger coefficients lead to a smoother surface, but excessively small coefficients may cause tetrahedron inversion. In our experiments, we choose $w_{1}=5\times 10^{-6},w_{2}=2\times 10^{-5}$ and apply a cosine increasing schedule.
@@ 6 Conclusion
We introduced TetSphere splatting, a Lagrangian geometry representation for high-quality 3D shape modeling. Our computational framework leverages geometric energy optimization and differentiable rendering to deform TetSpheres into complex shapes, producing results that demonstrate superior mesh quality compared to state-of-the-art methods.
Limitations. TetSphere splatting has limitations, including the generation of piecewise meshes rather than a globally consistent tetrahedral mesh and the lack of strong topological guarantees. More details are discussed in Appendix K. Future work could extend TetSphere splatting to leverage direct 3D supervision with volumetric data, eliminating the need for rendering to images. The current TetSphere splatting cannot guarantee topology preservation due to the union of all tetrahedral spheres. This highlights the need for future development in shape generation that can better adhere to topology constraints.
Appendix A Comparison with Single-Primitive Approaches in Inverse Rendering
Figure 8: Qualitative comparisons with single-primitive approaches in inverse rendering (Nicolet et al., 2021; Palfinger, 2022).
Our method is related to single-primitive inverse rendering methods, such as  Nicolet et al. (2021) and Palfinger (2022). While these existing methods use a single surface sphere for reconstruction, our method leverages a volumetric representation composed of multiple tetrahedral spheres instead of just one. Both Nicolet et al. (2021); Palfinger (2022) and our approach employ regularizations to prevent extreme deformations that can compromise geometric quality, thereby inherently limiting each sphere’s expressivity to a certain extent. Nicolet et al. (2021); Palfinger (2022) address this limitation by using intermediate remeshing to accommodate topological changes and manage drastic deformations when the target shape significantly differs from the sphere. However, remeshing results in the reparameterization of the texture, which poses challenges in the context of 3D shape generation pipelines. Maintaining a consistent texture parameterization throughout the optimization process is crucial, as the texture image itself is an optimization variable (see Appendix H). Additionally, remeshing can introduce undesired meshing complications when the surface undergoes topological changes. To prevent topological changes during each iteration, a single primitive must be initialized with a topology that matches the target. By contrast, our method eliminates the need for intermediate remeshing by utilizing multiple TetSpheres alongside stronger regularization terms, including biharmonic energy and non-inversion constraints. By leveraging multiple TetSpheres, each TetSphere undergo a smaller deformations, ensuring robust shape reconstruction.
Table 4 and Fig. 8 illustrate the quantitative and qualitative comparisons of our method against  Nicolet et al. (2021) and Palfinger (2022). We used their publicly available codebases and performed multi-view reconstruction on our dataset as described in Sec. 5.2. Our method outperforms these approaches, particularly for shapes with complex topologies. Both  Nicolet et al. (2021) and Palfinger (2022) use a single sphere for reconstruction, which limits their ability to handle objects with multiple holes, such as the dress and the sorter. For the dress example (the second and the third columns in Fig. 8), both baseline methods exhibit unrealistic closure of the open areas. In contrast, our approach accurately represents the thin regions and preserves the open areas due to the use of multiple TetSpheres. For the sorter example (the fourth and the fifth columns), the results from Nicolet et al. (2021) exhibit folded, overlapping surfaces with noticeable crumpled regions. This occurs because their deformation regularization is only on the surface and cannot effectively penalize large bending. Our tetrahedral-based method, with its volumetric regularization, mitigates these issues.
Table 4: Quantitative comparisons with single-primitive approaches in inverse rendering (Nicolet et al., 2021; Palfinger, 2022). We omit MR and CC Diff., as all three methods achieve identical values.
Method Cham. $\downarrow$ Vol. IoU $\uparrow$ ALR $\uparrow$
Nicolet et al. (2021) 0.0213 0.6184 0.5235
Palfinger (2022) 0.0191 0.6583 0.6332
Ours 0.0184 0.6844 0.6602
Appendix B Additional Ablation Studies
B.1 Ablation Study on the Number of TetSpheres
Figure 9: Qualitative reconstruction quality comparison using different numbers of TetSpheres. At one extreme, using only a single TetSphere leads to poor reconstruction, unable to capture the intricate topology of the target shapes. As the number of TetSpheres increases, the reconstruction quality improves. Beyond a certain point, however, the quality stabilizes although the tessellation density becomes higher. Our chosen number of TetSpheres ($323$ and $127$) strikes a balance between achieving high reconstruction quality and maintaining reasonable computational efficiency.
The number of TetSpheres in our method is determined by the initialization algorithm described in Sec. 4 and Appendix G, controlled by the scaling and offset parameters $\alpha$ and $\beta$ that define the initial radius of each TetSphere. Intuitively, larger values of $\alpha$ and $\beta$ result in a sparser distribution of TetSpheres. In this section, we present an ablation study on the reconstruction performance with respect to different numbers of TetSpheres.
We select two examples from the multi-view reconstruction experiment, the dress and the sorter, as these represent the most challenging cases due to their complex topology. For each example, we evaluate five different combinations of $\alpha$ and $\beta$, resulting in {$1,59,87,323,434$} TetSpheres for the dress and {$1,31,69,127,239$} for the sorter. Notably, $323$ and $127$ TetSpheres are the numbers used in our primary method.
The reconstruction results are illustrated in Fig. 9, with corresponding metrics for Chamfer distance, volume IoU, and ALR shown in Fig. 10(a). At one extreme, using only a single TetSphere leads to poor reconstruction, unable to capture the intricate topology of the target shapes. As the number of TetSpheres increases, the reconstruction quality improves significantly. Beyond a certain point, however, such as with $434$ and $239$ TetSpheres, the quality stabilizes, and the metrics indicate minimal further improvement – although the tessellation density becomes higher.
An excessive number of TetSpheres increases computational costs due to the additional memory required to store the vertex positions. Our chosen number of TetSpheres strikes a balance between achieving high reconstruction quality and maintaining reasonable computational efficiency.
B.2 Ablation Study on the Number of Tetrahedra per TetSphere
Figure 10: (a) Quantitative comparison using different numbers of TetSpheres. As the number of TetSpheres increases, both the reconstruction accuracy (Chamfer distance and Vol. IoU) and the surface quality (ALR) improve. Beyond $434$ and $239$, however, the metrics indicate minimal further improvement. (b) Ablation study on the trade-off between the average number of tetrahedra per TetSphere and the total number of TetSpheres. For each setting of the number of TetSpheres, increasing the number of tetrahedra per TetSphere improves reconstruction accuracy. However, the improvements are less pronounced compared to increasing the number of TetSpheres.
We conducted an ablation study to analyze the trade-off between the average number of tetrahedra per TetSphere and the total number of TetSpheres required to represent the surface. We used four sets of $\alpha$ and $\beta$ values to obtain different granularities of TetSphere counts, averaging {$9$, $38$, $110$, $213$} across all examples in the multi-view reconstruction experiment. For each granularity, we varied the average number of tetrahedra per TetSphere, testing five configurations: {6000, 9000, 12000, 15000, 18000}. The Chamfer Distance results for these tests are shown in Fig. 10(b).
For each granularity setting, increasing the number of tetrahedra per TetSphere improved reconstruction accuracy. However, the improvements are less significant compared to increasing the number of TetSpheres. While more tetrahedra per TetSphere can provide finer detail within a single TetSphere, the primary determinant of reconstruction accuracy is the overall number of TetSpheres. This is because distributing the surface representation across a larger number of TetSpheres allows for better coverage and adaptability to complex geometric features. Therefore, the number of tetrahedra within each TetSphere plays a secondary role in the overall surface reconstruction quality.
Appendix C Formulations of Metrics
Area-Length Ratio (ALR): The ALR is defined as following the definition of TriangleQ in Frey & Borouchaki (1999); Xu et al. (2024a):
$\text{ALR}=\frac{1}{N}\sum_{i=1}^{N}\frac{6}{\sqrt{3}}\frac{A_{i}}{P_{i}h_{t}},$
where $N$ is the total number of triangles in the mesh, $A_{i}$ is the triangle area, $P_{i}$ is the half-perimeter, and $h_{t}$ is the longest edge length. This metric ranges from 00 to $1$, with higher values indicating that the mesh consists primarily of equilateral triangles, thus reflecting superior triangle quality.
Manifoldness Rate (MR): The MR checks whether the mesh is a closed manifold. The MR is reported as the percentage of meshes in the evaluation dataset that qualify as closed manifolds:
MR=Number of manifold meshesTotal number of $\text{MR}=\frac{\text{Number of manifold meshes}}{\text{Total number of meshes}}\times 100\%.$
Connected Component Discrepancy (CC Diff.): This metric measures the difference in the number of connected components between the reconstructed mesh and the ground-truth shape. Let $C^{(i)}_{\text{rec}}$ and $C^{(i)}_{\text{gt}}$ denote the number of connected components in the reconstructed and ground-truth meshes, respectively. The CC Diff. is defined as:
[…]
which helps identify floating artifacts or discontinuities, indicating the structural integrity and cohesion of the reconstructed shape.
Appendix D Simple Study on Tetrahedron Inversion
In the inset figure (left), we provide a 2D illustration showing how inversion can occur, with the inverted region highlighted in red. To further illustrate the importance of our regularization in preventing tetrahedron inversion, we conducted a simple study using a target shape of a sphere that is smaller than the initial TetSphere. As depicted in the inset figure (right), without regularization, the rendering loss alone pushes the surface vertices of the initial TetSphere directly onto the target sphere’s surface. This results in a significant number of inverted tetrahedra, as highlighted in red. In contrast, with our volumetric regularization in place, the reconstructed shape successfully adheres to the target surface without any inverted tetrahedra, ensuring a high-quality reconstruction.
Figure 11: Qualitative results on single-view reconstruction. We use Wonder3d (Long et al., 2023) to generate six multi-view images with predefined camera poses. The optimization objective includes rendering loss $l_{1}$ norm on tone-mapped color and MSE on the alpha mask) and normal loss (cosine loss on normals), following Munkberg et al. (2022). More details are described in Sec. 5.1
Table 5: Additional results on multi-view reconstruction.
Metric NIE FlexiCubes 2DGS DMesh Ours
F-Score $\uparrow$ 0.486 0.502 0.632 0.605 0.653
Normal Consis.$\uparrow$ 0.718 0.734 0.812 0.745 0.839
Edge Cham. $\downarrow$ 0.039 0.034 0.015 0.049 0.014
Edge F-Score $\uparrow$ 0.016 0.196 0.219 0.193 0.269
Time (sec.) $\downarrow$ 6436 57.67 691 1434 934
# Triangles 151,223.6 30,285.7 270,235.5 25,767.7 70,616.0
Appendix E Additional Results
E.1 More Quantitative Results on Multi-view Reconstruction
Table 5 shows comparisons of multi-view reconstruction on more evaluation metrics.
E.2 Qualitative Results on Single-view Reconstruction
@@
Fig. 11 shows the results of single-view reconstruction performed on the GSO dataset. Our TetSphere splatting demonstrates superior mesh quality, effectively capturing sharp geometric features, including the boundaries of the shoes.
Figure 12: Qualitative comparison on image-to-3D generation with surface mesh visualizations and rendered normal maps. Our generated meshes exhibit less bumpiness and high mesh quality with more regular triangle meshing. We use multi-view images generated using the coarse NeRF fitting stage of DreamCraft3D, sampled on a Fibonacci sphere, and optimize a $2048\times 2048$ texture image directly using rendering loss. More details are described in Appendix I.
Figure 13: More results on text-to-3D shape generation. TetSphere splatting is optimized using rendering and SDS losses. The SDS loss for geometry is based on the Normal-depth diffusion model in Qiu et al. (2023), and loss for texture uses the albedo diffusion model from the same work. More details are described in Appendix I.
E.3 Qualitative Results on Image-to-3D Generation
Fig. 12 shows qualitative comparisons of image-to-3D generation with surface mesh visualizations and rendered normal maps. Our generated meshes exhibit minimal noise and high quality, featuring more regular triangle meshing.
E.4 More Results on Text-to-3D Shape Generation
@@
Appendix F Tetrahedron and its Deformation Gradient
Following (Sifakis & Barbic, 2012), we treat a tetrahedron as a piecewise linear element. The initial (undeformed) positions of the four vertices of a tetrahedron are denoted by $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\mathbf{X}^{(3)},\mathbf{X}^{(4)}]$, with each $\mathbf{X}^{(i)}\in\mathbb{R}^{3}$. Similarly, the positions of the four vertices of a deformed tetrahedron are represented by $\mathbf{x}=[\mathbf{x}^{(1)},\mathbf{x}^{(2)},\mathbf{x}^{(3)},\mathbf{x}^{(4)}]$, where each $\mathbf{x}^{(i)}\in\mathbb{R}^{3}$. The deformation gradient $\mathbf{F}\in\mathbb{R}^{3\times 3}$, which quantifies the local deformation of the tetrahedron, is given by:
@@ Appendix F Tetrahedron and its Deformation Gradient
Appendix G Formulation of TetSphere Initialization
@@
Appendix H Texture and PBR Material Optimization
Material optimization is facilitated by using differentiable rasterizers (Laine et al., 2020), which adjust the textures or materials to closely match the input multi-view color images. A significant advantage of TetSphere splatting is that the deformation of tetrahedral spheres does not alter the surface topology. Unlike methods such as DMTet, which require isosurface extraction and subsequent texture parameterization at each step due to potential changes in the underlying shape, our method necessitates only a single texture parameterization at the beginning of optimization. This parameterization remains consistent throughout the process, significantly enhancing the efficiency of texture optimization.
For scenarios with dense input views, we have found that using textured images as optimization variables is straightforward and yields high-quality results. In cases with sparse input views, we adopt a two-layer MLP that takes the surface vertex positions as inputs and outputs the material parameters, a practice in line with existing methods (Sun et al., 2023; Qiu et al., 2023).
Appendix I More Implementation Details
When a surface mesh is required, we extract the surface triangles from each TetSphere to generate surface spheres, and then perform a union of all these surface spheres to obtain the final surface mesh. While there is no theoretical guarantee of manifoldness in the resulting mesh from union operation, our TetSphere optimization is highly regularized to maintain geometric quality. As a result, we did not encounter non-manifold issues in our experiments.
For image-to-3D shape generation, we obtain multi-view images from the initial stage of DreamCraft3D (coarse NeRF fitting only), generating $360$ views sampled on a Fibonacci sphere (González, 2010). The texture optimization directly employs a $2048\times 2048$ 2D texture image, circumventing the need for additional neural networks. Here, the optimization objective $\mathbf{\Phi}(\cdot)$ focuses solely on the rendering loss. For text-to-3D generation, we use the initial stage of RichDreamer (Qiu et al., 2023) to obtain multi-view images. We optimize our TetSphere splatting using both the rendering loss and the SDS loss. The SDS loss for geometry is calculated using the Normal-depth diffusion model as described in Qiu et al. (2023). The SDS loss for texture leverages the albedo diffusion model from the same work.
Regarding baselines: For image-to-3D generation, we compare TetSphere splatting with Magic123 and Dreamcraft3D (Sun et al., 2023). Both are multi-stage methods starting with NeRF optimization followed by DMTet to optimize mesh and texture. For text-to-3D, our comparison focuses on RichDreamer, a state-of-the-art method known for incorporating PBR materials and generating shapes, showcasing notable results in 3D generation from text prompts.
Appendix J Related work – Text-to-3D content generation